OpenStudy (anonymous):

A problem in nine chapter on the mathematical art can be represented by this system of equations. Solve the system 9x-y=11 6x-y=-16

OpenStudy (anonymous):

or, you could look at it as: 9x-y=11 y=16+6x

OpenStudy (anonymous):

then sub equation 2 into equation 1, : )

OpenStudy (anonymous):

I hate math

OpenStudy (anonymous):

it certainly isn't the easiest subject in the world, lol

OpenStudy (anonymous):

im failing in my class

OpenStudy (anonymous):

I'm sorry to hear that; that isn't the nicest feeling in the world is it... although, I'm sure though that if I could find a way to pass it, I you will too, lol

OpenStudy (anonymous):

haha im never gonna get math

OpenStudy (anonymous):

I don't believe that for a second, lol

OpenStudy (anonymous):

I failed 2 years of prealgebra and barely made it past algebra

OpenStudy (anonymous):

I still don't believe that it's impossible for you to get it - it may feel absolutely impossibly difficult, but I believe you can get it eventually... if you're not too hard on yourself : )

OpenStudy (anonymous):

too late I already have it set in my mind that im a failure

OpenStudy (anonymous):

You're here; and that's more than where any failure would be.

OpenStudy (anonymous):

cause im trying to do something so I don't get sent to a girls town

OpenStudy (anonymous):

Do you know how to continue this problem?

OpenStudy (anonymous):

by substituting the second equation into the first, you get: 9x - (16+6x) = 11

OpenStudy (anonymous):

which is equation to: 9x -16 -6x = 11

OpenStudy (anonymous):

can you solve for x?

OpenStudy (anonymous):

here's a good question; can you cook?

OpenStudy (anonymous):

... nvm, lol. simplifying for x gives you: 3x=27 x=9 substitute x=9 into the original equation to get: y=-11/81

OpenStudy (anonymous):

And your not a failure; this stuff is tough for a lot of people.... All the best!